OpenStudy (anonymous):

integral sqrt (x^2+64)/ 5x^2 please help

zepdrix (zepdrix):

\(\Large\bf \color{#008353}{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits \frac{\sqrt{x^2+64}}{5x^2}dx\]

zepdrix (zepdrix):

Ummm so you have a few different options. A `Trig Sub` is one of them. Are you familiar with that method?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\Large\bf\sf x=8\tan \theta\]Do you understand why I chose that for the substitution?

OpenStudy (anonymous):

yes becuase the triangle in my book has us do that i also know that \[dx =8\sec ^2\]

zepdrix (zepdrix):

\[\Large\bf\sf dx=8\sec^2\theta \;d \theta\]Ok good.

zepdrix (zepdrix):

Plug all the stuff in! :)

OpenStudy (anonymous):

thats where I'm getting stuck

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits\limits \frac{\sqrt{x^2+64}}{5x^2}dx\quad=\quad \int\limits\limits \frac{\sqrt{64\tan^2\theta+64}}{5\cdot64\tan^2\theta}(8\sec^2\theta\;d \theta)\]

zepdrix (zepdrix):

That look ok so far? Simplifying will be a little tricky :)

OpenStudy (anonymous):

\[I=\int\limits \frac{ \sqrt{x^2+64} }{ 5x^2 }dx\] \[=\int\limits \frac{ \left( x^2+64 \right) }{5x^2\sqrt{x^2+64} }dx\] \[Put~ x=\frac{ 1 }{ t },dx=-\frac{ 1 }{ t^2 }dt\] \[I=\int\limits \frac{ \frac{ 1 }{ t^2 }+64 }{ \frac{ 5 }{ t^2 }\sqrt{\frac{ 1 }{ t^2 }+64} }\left( \frac{ -1 }{ t^2 }\right) dt\] \[=\frac{ 1 }{5 }\int\limits \frac{ -1\left( 1+64t^2 \right) }{t \sqrt{1+64t^2} }dt\] \[=-\frac{ 1 }{5 }\int\limits \frac{ \left( 1+64t^2 \right)t~dt }{ t^2\sqrt{1+64t^2} }\] put \[\sqrt{1+64t^2}=y,~64 ~t^2=y^2-1,t^2=\frac{ y^2-1 }{64 },2t~dt=\frac{ 2ydy }{64 }\] \[tdt=\frac{ ydy }{64 }\] \[I=-\frac{ 1 }{5}\int\limits \frac{ y^2*\frac{ y~dy }{ 64 } }{\frac{ y^2-1 }{64 }y }=-\frac{ 1 }{5 }\int\limits \frac{ y^2dy }{y^2-1 }\] \[=-\frac{ 1 }{5 }\int\limits \frac{ y^2-1+1 }{ y^2-1 }dy=-\frac{ 1 }{5 }\left[ \int\limits dy+\int\limits \frac{ 1 }{ (y+1)(y-1) }dy \right]\] complete it.

OpenStudy (anonymous):

make partial fractions and complete.

OpenStudy (anonymous):

Also consider doing integration by parts \[ \text{dv}=\frac{1}{5 x^2}\\ u=\sqrt{x^2+64}\\ v=-\frac 1 {5 x}\\ du=\frac{x}{\sqrt{x^2+64}}\\ Integral= -\frac{\sqrt{x^2+64}}{5 x}+ \int \frac 1 {5 x} \frac{x}{\sqrt{x^2+64}}dx=\\ -\frac{\sqrt{x^2+64}}{5 x}+ \int \frac 1 {5 } \frac{ dx}{\sqrt{x^2+64}} =\\ \frac{1}{5} \left(-\frac{\sqrt{x^2+64} }{x}+\sinh ^{-1}\left(\frac{x}{8}\right)\right) +C \]

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

\[\frac{ 1 }{ \left( y+1 \right)\left( y-1 \right) }=\frac{ A }{y+1}+\frac{ B }{y-1 }\] 1=A(y-1)+B(y+1) put y=1 1=A(1-1)+B(1+1) B=1/2 put x=-1 1=A(-1-1)+B(-1+1) A=-1/2 \[\int\limits \frac{ 1 }{\left( y+1 \right)\left( y-1 \right) }dy=\frac{ 1 }{2 } \int\limits \left( \frac{ -1 }{ y+1 }+\frac{ 1 }{y-1 } \right)dy\] \[=\frac{ 1 }{ 2 }\left( -\ln \left( y+1 \right)+\ln \left( y-1 \right) \right)=-\frac{ 1 }{2 }\ln \frac{ y+1 }{ y-1 }\] \[I=-\frac{ 1 }{ 5 }\left[ y-\frac{ 1 }{ 2 }\ln \frac{ y+1 }{ y-1 } \right]+c\] \[I=-\frac{ 1 }{5 }\left[ \sqrt{1+64t^2}-\frac{ 1 }{2 }\ln \frac{ \sqrt{1+64t^2}+1 }{\sqrt{1+64t^2}-1 } \right]+c\] \[I=\frac{ -1 }{5 }\left[ \sqrt{1+\frac{ 64 }{ x^2 }}-\frac{ 1 }{ 2 }\ln \frac{ \sqrt{1+\frac{ 64 }{ x^2 }}+1 }{\sqrt{1+\frac{ 64 }{ x^2 }}-1 } \right]+c\] \[I=-\frac{ 1 }{5 }\left[ \frac{ \sqrt{x^2+64} }{ x }-\frac{ 1 }{ 2 } \ln\frac{\sqrt{x^2+64} +x }{ \sqrt{x^2+64}-x }\right]+c\]

OpenStudy (anonymous):

as suggested by Mr. eliassaab \[\int\limits \frac{ 1 }{ \sqrt{x^2+64} }dx,\ put~x=8\tan t,dx=8\sec ^2t~dt\] \[\int\limits \frac{ 8~\sec ^2 ~t~dt }{ 8\sqrt{\tan ^2t+1} }=\int\limits \frac{ \sec ^2t~dt }{ \sec t }=\int\limits \sec t~dt=\ln \left| \sec t+\tan t \right|\] \[=\ln \left| \tan t+\sqrt{\tan ^2 t+1} \right|=\ln \left| \frac{ x }{ 8 }+\frac{ \sqrt{x^2+64} }{ 8 } \right|\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }\ln \frac{ \sqrt{x^2+64}+x }{ \sqrt{x^2+64}-x }\times \frac{ \sqrt{x^2+64}+x }{ \sqrt{x^2+64}+x }\] \[=\frac{ 1 }{ 2 } \ln \frac{ \left( \sqrt{x^2+64}+x \right)^2 }{ x^2+64-x^2 }\] \[=\frac{ 1 }{ 2 }\ln \left( \frac{ \sqrt{x^2+64}+x }{ 8 } \right)^2\] \[=2\times \frac{ 1 }{ 2 }\ln \left( \frac{ \sqrt{x^2+64}+x }{ 8 } \right)=\ln \left( \frac{ \sqrt{x^2+64}+x }{ 8 } \right)\]